The polypeptide chain of the chaperone protein GroEL from Escherichia colihas a molecular mass of 57200 Da and does not contain tryptophan or cystines. The A280of a 1 mg/mL solution of the purified protein in a 1 cm pathlength cuvette is 0.160. calculate the number of Tyr in the polypeptide chain.
Answer #1

Protein Concentration in mg/ml = A280*Dilution Factor*molecular weight/Extinction coefficient

1= 0.160*57200/Extinction coefficient

  Extinction coefficient= 9152

for one residue of tyrosine the Extinction coefficient is 1280 so if the value is 9152 so the number of residue will be

= 9152/1280= 7.15 residues